# Schrödinger as a Quantum Programmer: Estimating Entanglement via Steering

### v3

@@ Acknowledgments
We are especially grateful to Gus Gutoski, for providing the main idea of the quantum interactive proof detailed in Figure 1, back in September 2013. We also thank Paul Alsing, Eric Chitambar, Zoe Holmes, and Wilfred Salmon for insightful discussions, and Ludovico Lami, Bartosz Regula, and Alexander Streltsov for the same, as well as pointing us to [72]. AP, SR, and MMW acknowledge support from the National Science Foundation under Grant No. 1907615.
@@ Appendix A Proof of Theorem 1
Proof of Theorem 1. Recall that an entanglement-breaking channel can be rewritten as
@@
After optimizing over every element of $\operatorname{EB}_{R\to A^{\prime}}$, which denotes the set of all entanglement-breaking channels with input system $R$ and output system $A^{\prime}$, and realizing that optimizing over measurements in $\mathcal{E}_{R\rightarrow A^{\prime}}$ induces a pure-state decomposition of $\rho_{AB}$ and optimizing over preparation channels in $\mathcal{E}_{R\rightarrow A^{\prime}}$ gives the spectral norm of $\psi_{A}^{x}$, we find the claimed formula for the acceptance probability, when combined with the development in Appendices B and C:
@@ Appendix B Alternative Proof of Equation (19)
This appendix provides an alternative proof for Theorem 1 in [61]. This proof relies on Uhlmann’s theorem [52], the triangle inequality, and the Cauchy–Schwarz inequality. See also [72, Lemma 1].
@@
The following formula holds
@@
where $\{(p(x),\psi_{AB}^{x})\}_{x}$ satisfies $\sum_{x}p(x)\psi_{AB}^{x}=\rho_{AB}$, all $\psi_{AB}^{x}$ are pure, and
@@ Appendix C Proof of Equation (20)
In this appendix, we show that the fidelity of separability of a bipartite state can be written in terms of the spectral norm, which was also observed in [61, Section 6.2].
@@
For a bipartite state, the following equality holds
@@ Appendix D Proof of Theorem 2
  • Maximization over the parameters of $U^{x}_{A}(\Theta^{x})$ is a maximization of $\langle\phi^{x}|\varphi^{x}_{A}|\phi^{x}\rangle$, which yields the value of $\left\|\varphi_{A}^{x}\right\|_{\infty}$.
@@
This proves that if the parameterized unitary circuits can express all possible unitary operators of their respective systems, the maximum acceptance probability equals $F_{s}(\rho_{AB})$. However, we note that any ansatz employed for the parameterized unitary circuits has limited expressibility. As such, the maximum acceptance probability obtained via the VQSA will also be closer to the actual value of $F_{s}(\rho_{AB})$ if we use a more expressive ansatz.
@@ Appendix E First Benchmarking SDP $\widetilde{F}_{s}^{1}$ and Proof of Equation (21)
We would ideally like to include a maximization over the set of all separable states, but it is well known to be computationally challenging to optimize over this set [3, 4]. Note that it is not generally possible to characterize the set of separable states using semi-definite constraints [74]. Instead, we can only approximate the set by constraining $\sigma_{AB}$ to have a positive partial transpose (PPT) [27, 28] and be $k$-extendible [29, 30], since all separable states satisfy these constraints. Let $\widetilde{F}_{s}^{1}(\rho_{AB})$ denote the resulting quantity, the square root of which is defined as follows:
$\sqrt{\widetilde{F}_{s}^{1}}(\rho_{AB},k)\coloneqq\max_{\begin{subarray}{c}X_{AB}\in\mathcal{L}(\mathcal{H}_{AB}),\\ \sigma_{AB^{k}}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Re}[\operatorname{Tr}[X_{AB}]]:\\ \begin{bmatrix}\rho_{AB}&X_{AB}\\ X_{AB}^{{\dagger}}&\sigma_{AB_{1}}\end{bmatrix}\geq 0,\\ \operatorname{Tr}[\sigma_{AB^{k}}]=1,\\ \sigma_{AB^{k}}=\mathcal{P}_{B^{k}}(\sigma_{AB^{k}}),\\ T_{B_{1\cdots j}}(\sigma_{AB_{1\cdots j}})\geq 0\quad\forall j\leq k\end{array}\right\},$ (63)
@@
We prove the inequalities in (21). Due to the containment discussed above, we note that
@@
It is a consequence of [67, Theorem II.7], the triangle inequality for sine distance [83], and the Fuchs-van-de-Graaf inequalities [84]. Indeed, consider that
@@
where EXT-PPTk denotes the set being optimized over in (63) and EXTk is the set of $k$-extendible states. Now recall that for all $\omega^{k}_{AB}\in\operatorname{EXT}_{k}$ [67, Theorem II.7’]
@@ Appendix F Second Benchmarking SDP $\widetilde{F}_{s}^{2}$ and Proof of Equation (23)
and $\mathcal{P}$ is a preparation channel. Observing that $\overline{\mathcal{P}}_{R\rightarrow A^{\prime}}$ is a $k$-extendible channel [77, 78, 79, 80], it follows that
$\max_{\mathcal{P}_{R\rightarrow A_{1}^{\prime}\cdots A_{k}^{\prime}}}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\overline{\mathcal{P}}_{R\rightarrow A^{\prime}}(\psi_{RAB})]$ $=\max_{\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in\text{EXT}_{k}}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\mathcal{E}_{R\rightarrow A^{\prime}}^{k}(\psi_{RAB})]$ (78)
$\eqqcolon\frac{1}{2}(1+\widetilde{F}_{s}^{2}(\rho_{AB},k)),$ (79)
@@
12($\frac{1}{2}(1+\widetilde{F}_{s}^{2}(\rho_{AB},k))=\max_{\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}\geq 0}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\operatorname{Tr}_{R}[T_{R}(\psi_{RAB})\Gamma^{\mathcal{E}^{k}}_{RA^{\prime}_{1}}]]:\\ \operatorname{Tr}_{A^{\prime k}}[\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}]=I_{R},\\ \Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}=\mathcal{P}_{A^{\prime k}}(\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}),\\ T_{A^{\prime}_{1\cdots j}}(\Gamma^{\mathcal{E}^{k}_{RA^{\prime k}}})\geq 0\quad\forall j\leq k\end{array}\right\},$ (80)
where $\Gamma^{\mathcal{E}^{k}}_{RA^{\prime}}$ is the Choi operator of $\mathcal{E}^{k}$, the map $T_{R}$ is the partial transpose map acting on system $R$, and $\mathcal{P}_{A^{\prime k}}$ is the channel that randomly permutes the systems $A^{\prime k}$.
@@
$\leq\max_{\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in\text{EXT}_{k}}\operatorname{Tr}[(\Pi_{A^{\prime}A}^{\operatorname{sym}}\otimes I_{RB})\mathcal{E}_{R\rightarrow A^{\prime}}^{k}(\psi_{RAB})]$ (83)
$=\frac{1+\widetilde{F}_{s}^{2}(\rho_{AB},k)}{2}.$ (84)
Consider the following bound for a $k$-extendible state $\omega_{AB}^{k}$ [67, Theorem II.7’]:
@@
We can use it and the result of [85, Lemma 7] to conclude that
@@
Then consider, for every fixed choice of $\mathcal{E}_{R\rightarrow A^{\prime}}^{k}$, there exists an entanglement-breaking channel $\mathcal{E}$ satisfying
@@
Since the inequality holds for every $\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in$ EXTk, it follows that
$\max_{\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in\text{EXT}_{k}}\operatorname{Tr}[(\Pi_{A^{\prime}A}^{\operatorname{sym}}\otimes I_{RB})\mathcal{E}_{R\rightarrow A^{\prime}}^{k}(\psi_{RAB})]\leq\frac{1+F_{s}(\rho_{AB})}{2}+\frac{2\left|R\right|\left|A^{\prime}\right|^{2}}{k}.$ (91)
@@ Appendix G Further Simulations and Details
For all the simulations in our work, the input states and the parameterized unitaries are generated using the hardware efficient ansatz (HEA) [86]. The HEA consists of several layers, where each layer consists of two parameters per qubit per layer, specifying rotations about the $x$- and $y$-axes. After each layer of rotations is a series of neighboring qubit CNOT gates. When using the HEA to generate the input states, we keep the rotation angles fixed, thus leading to a fixed input state. For the parameterized unitaries, the rotation angles are parameters and are optimized over.
In Figure 10(a), we report simulation results after generating a random bipartite product state, with each partition containing two qubits. To guarantee a product state, we remove all the CNOT gates from the HEA that generates the input state $\rho$. We calculated the fidelity of separability using both the local reward function of the VQSA and the benchmark $\widetilde{F}^{1}_{s}$, the latter discussed in Appendix E.
@@
We generated all parameterized unitary circuits in the following fashion. We used the Qiskit Aer simulator and Qiskit’s Simultaneous Perturbation Stochastic Approximation (SPSA) optimizer to perform the classical optimization. All other details can be found in Table 1. The local reward function of the VQSA requires more classical processing (like picking a qubit at random to measure) and seems to require more iterations to reach the right value. However, these downsides are outweighed by the fact that it is less susceptible to the emergence of barren plateaus. More details about the local cost function can be found in Appendix L.
@@ Appendix I Multipartite Scenarios
In this appendix, we discuss a multipartite generalization of mixed states’ separability tests.
@@
$\rho_{A_{1}\cdots A_{M}}=\sum_{x\in\chi}p(x)\psi^{x,1}_{A_{1}}\otimes\cdots\otimes\psi^{x,M}_{A_{M}},$ (92)
where $\psi_{A_{i}}^{x,i}$ is a pure state for every $x\in\mathcal{X}$ and $i\in\{1,\ldots,M\}$.
@@
$\max_{\sigma_{A_{1}\cdots A_{M}}\in M-\operatorname{SEP}}F(\rho_{A_{1}\cdots A_{M}},\sigma_{A_{1}\cdots A_{M}})=\max_{\{(q(x),\varphi_{A_{1}\cdots A_{M}}^{x})\}_{x}}\sum_{x}q(x)F_{s}(\varphi^{x}_{A_{1}\cdots A_{M}}),$ (93)
@@
$\max_{\mathcal{E}\in\operatorname{EB}_{M-1}}\operatorname{Tr}[\Pi_{(A^{M-1})^{\prime}(A^{M-1})}^{\operatorname{sym}}\mathcal{E}_{R\rightarrow A^{\prime}_{1}\cdots A^{\prime}_{M-1}}(\psi_{RA^{M}})]=\frac{1}{2}\left(1+\max_{\sigma_{A_{1}\cdots A_{M}}\in M-\operatorname{SEP}}F(\rho_{A_{1}\cdots A_{M}},\sigma_{A_{1}\cdots A_{M}})\right).$ (97)
@@
where $\{\mu^{x}_{R}\}_{x}$ is a rank-one POVM and $\{\phi^{x,i}_{A^{\prime}_{i}}\}_{x,i}$ is a set of pure states. The prover sends the systems $(A^{M-1})^{\prime}=A^{\prime}_{1}\cdots A^{\prime}_{M-1}$ to the verifier. Now, the verifier performs a collective swap test on $A_{1}\cdots A_{M}$, as depicted at the final part of the circuit diagram in Figure 11. The acceptance probability of this interactive proof system is thus given by
@@
is the projector onto the symmetric subspace of $A^{\prime}$ and $A$ and $F_{(A_{1}\cdots A_{M-1})^{\prime}A_{1}\cdots A_{M-1}}$ is a tensor product of individual swaps $F_{A^{\prime}_{i}A_{i}}$, for $i\in\{1,\ldots,M\}$. That is,
$F_{(A_{1}\cdots A_{M-1})^{\prime}A_{1}\cdots A_{M-1}}=\bigotimes_{i=1}^{n}F_{A^{\prime}_{i}A_{i}}.$ (101)
@@
where $\mathcal{V}_{R^{\prime}S\rightarrow DG}$ is the unitary channel corresponding to the unitary operator $V_{R^{\prime}S\rightarrow DG}$. By the reasoning similar to that in (25), (32), and (33), we find that
@@
where we have used the shorthand $\mathcal{V}\equiv\mathcal{V}_{R^{\prime}S\rightarrow DG}$. Consider that
@@
$\mathcal{W}_{G\rightarrow R^{\prime}S}(\cdot)\coloneqq(V_{R^{\prime}S\rightarrow DG})^{{\dagger}}(|1\rangle\!\langle 1|_{D}\otimes(\cdot)_{G})V_{R^{\prime}S\rightarrow DG}.$ (124)
Then, the acceptance probability is given by
@@
Observe that the map $\rho_{AB}\mapsto\Pi^{\operatorname{sym}}_{AA^{\prime}}(\rho_{AB}\otimes I_{A^{\prime}})\Pi^{\operatorname{sym}}_{AA^{\prime}}$ is proportional to that used in a $1\to 2$ universal cloning machine [93, Eq. (17)]. If $\rho_{AB}$ is pure, so that we write it as $\psi_{AB}$, then the following inequality holds:
@@
Quantum statistical zero-knowledge (QSZK) consists of all problems that can be solved by the interaction between a quantum verifier and a quantum prover, such that the verifier accumulates statistical evidence about the answer to a decision, but does not learn anything other than the answer by interacting with the prover [94, 56]. A complete problem for this class is quantum state distinguishability, in which the goal is to decide whether two states $\rho_{0}$ and $\rho_{1}$, generated by quantum circuits, are far or close in trace distance [94]. This is a nice problem for understanding the basics of the QSZK complexity class: the interaction begins with the verifier picking one of the states uniformly at random, recording the choice as a bit $x$, and then sending the chosen state $\rho_{x}$ to the prover over a quantum channel. The prover can then perform the optimal Helstrom measurement [95, 96] to distinguish the states which has success probability equal to
@@
Let us define the local and global reward functions. Let $Z_{i}$ be the event of measuring zero in the $i$th register. We then set the local reward function to be the probability of measuring zero in a register chosen uniformly at random; that is, it is given by the following:
@@
We are interested in determining inequalities related to the global and local reward functions, and the following analysis employs the same ideas used in [81, Appendix C]. Using DeMorgan’s laws, we find that
@@
$\Pr\!\left(\bigcap_{i}Z_{i}\right)$ $\leq 1-\frac{1}{n}\sum_{i}\left(1-\Pr\!\left(Z_{i}\right)\right)$ (174)
$=\frac{1}{n}\sum_{i}\Pr\!\left(Z_{i}\right).$ (175)

### v4

@@ Acknowledgments
We are especially grateful to Gus Gutoski for providing the main idea of the quantum interactive proof detailed in Figure 1, back in September 2013. We also thank Paul Alsing, Eric Chitambar, Zoë Holmes, and Wilfred Salmon for insightful discussions, and Ludovico Lami, Bartosz Regula, and Alexander Streltsov for the same, as well as pointing us to [73]. AP, SR, and MMW acknowledge support from the National Science Foundation under Grant No. 1907615 and from AFRL under agreement no. FA8750-23-2-0031.
This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes, notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government.
@@ Appendix A Proof of Theorem 1
Proof of Theorem 1. Recall that an entanglement-breaking channel can be rewritten as [51]
@@
After optimizing over every element of $\operatorname{EB}_{R\to A^{\prime}}$, which denotes the set of all entanglement-breaking channels with input system $R$ and output system $A^{\prime}$, and realizing that optimizing over measurements in $\mathcal{E}_{R\rightarrow A^{\prime}}$ induces a pure-state decomposition of $\rho_{AB}$ and optimizing over preparation channels in $\mathcal{E}_{R\rightarrow A^{\prime}}$ gives the spectral norm of $\psi_{A}^{x}$, we arrive at the claimed formula for the acceptance probability, when combined with the development in Appendices B and C:
@@ Appendix B Alternative Proof of Equation (19)
This appendix provides an alternative proof for Theorem 1 in [62]. This proof relies on Uhlmann’s theorem [52], the triangle inequality, and the Cauchy–Schwarz inequality. See also [73, Lemma 1].
@@
For a state $\rho_{AB}$, the following formula holds
@@
where the pure-state ensemble $\{(p(x),\psi_{AB}^{x})\}_{x}$ satisfies $\sum_{x}p(x)\psi_{AB}^{x}=\rho_{AB}$, all $\psi_{AB}^{x}$ are pure, and
@@ Appendix C Proof of Equation (20)
In this appendix, we show that the fidelity of separability of a bipartite state can be written in terms of the spectral norm, which was also observed in [62, Section 6.2].
@@
For a bipartite state $\rho_{AB}$, the following equality holds
@@ Appendix D Proof of Theorem 2
  • Maximization over the parameters of $U^{x}_{A}(\Theta^{x})$ is a maximization of $\langle\phi^{x}|\varphi^{x}_{A}|\phi^{x}\rangle$, which yields the value $\left\|\varphi_{A}^{x}\right\|_{\infty}$.
@@
This proves that the maximum acceptance probability equals $F_{s}(\rho_{AB})$ if the parameterized unitary circuits express all possible unitary operators acting on their respective systems. However, we note that any ansatz employed for the parameterized unitary circuits has limited expressibility. As such, the maximum acceptance probability obtained via the VQSA, in principle, will also be closer to the actual value of $F_{s}(\rho_{AB})$ if we use a more expressive ansatz.
@@ Appendix E First Benchmarking SDP $\widetilde{F}_{s}^{1}$ and Proof of Equation (21)
We would ideally like to include a maximization over the set of all separable states, but it is well known to be computationally challenging to optimize over this set [3, 4]. Furthermore, it is not generally possible to characterize the set of separable states using semi-definite constraints [75]. Instead, we approximate the set by constraining $\sigma_{AB}$ to have a positive partial transpose (PPT) [27, 28] and be $k$-extendible [29, 30], since all separable states satisfy these constraints. Let $\widetilde{F}_{s}^{1}(\rho_{AB})$ denote the resulting quantity, the square root of which is defined as follows:
$\sqrt{\widetilde{F}_{s}^{1}}(\rho_{AB},k)\coloneqq\max_{\begin{subarray}{c}X_{AB}\in\mathcal{L}(\mathcal{H}_{AB}),\\ \sigma_{AB^{k}}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\operatorname{Re}[\operatorname{Tr}[X_{AB}]]:\\ \begin{bmatrix}\rho_{AB}&X_{AB}\\ X_{AB}^{{\dagger}}&\sigma_{AB_{1}}\end{bmatrix}\geq 0,\\ \operatorname{Tr}[\sigma_{AB^{k}}]=1,\\ \sigma_{AB^{k}}=\mathcal{P}_{B^{k}}(\sigma_{AB^{k}}),\\ T_{B_{1\cdots j}}(\sigma_{AB_{1\cdots j}})\geq 0\quad\forall j\in\{1,\ldots,k\}\end{array}\right\},$ (63)
@@
We now prove the inequalities in (21). Due to the containment discussed above, note that
@@
It is a consequence of [68, Theorem II.7’], the triangle inequality for sine distance [84], and the Fuchs-van-de-Graaf inequalities [85]. Indeed, consider that
@@
where EXT-PPTk denotes the set of $\sigma_{AB}$ being optimized over in (63) and EXTk is the set of $k$-extendible states. Now recall that for all $\omega^{k}_{AB}\in\operatorname{EXT}_{k}$ (see [68, Theorem II.7’] and also just above [68, Theorem II.2] for their norm convention)
@@ Appendix F Second Benchmarking SDP $\widetilde{F}_{s}^{2}$ and Proof of Equation (23)
and $\mathcal{P}$ is an arbitrary channel. Observing that $\overline{\mathcal{P}}_{R\rightarrow A^{\prime}}$ is a $k$-extendible channel [78, 79, 80, 81], it follows that
$\max_{\mathcal{P}_{R\rightarrow A_{1}^{\prime}\cdots A_{k}^{\prime}}}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\overline{\mathcal{P}}_{R\rightarrow A^{\prime}}(\psi_{RAB})]$ $=\max_{\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in\text{EXT}_{k}}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\mathcal{E}_{R\rightarrow A^{\prime}}^{k}(\psi_{RAB})],$ (78)
@@
$\max_{\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}\geq 0}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\operatorname{Tr}_{R}[T_{R}(\psi_{RAB})\Gamma^{\mathcal{E}^{k}}_{RA^{\prime}_{1}}]]:\\ \operatorname{Tr}_{A^{\prime k}}[\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}]=I_{R},\\ \Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}=\mathcal{P}_{A^{\prime k}}(\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}})\end{array}\right\},$ (79)
where $\Gamma^{\mathcal{E}^{k}}_{RA^{\prime}}$ is the Choi operator of $\mathcal{E}^{k}$ and $\mathcal{P}_{A^{\prime k}}$ is the channel that randomly permutes the systems $A^{\prime k}$. We can add further PPT constraints to this SDP, which is still an upper bound on (4) and leads to our second benchmark $\widetilde{F}_{s}^{2}(\rho_{AB},k)$:
12($\frac{1}{2}(1+\widetilde{F}_{s}^{2}(\rho_{AB},k))\coloneqq\max_{\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}\geq 0}\left\{\begin{array}[c]{c}\operatorname{Tr}[\Pi_{A^{\prime}A}^{\operatorname{sym}}\operatorname{Tr}_{R}[T_{R}(\psi_{RAB})\Gamma^{\mathcal{E}^{k}}_{RA^{\prime}_{1}}]]:\\ \operatorname{Tr}_{A^{\prime k}}[\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}]=I_{R},\\ \Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}=\mathcal{P}_{A^{\prime k}}(\Gamma^{\mathcal{E}^{k}}_{RA^{\prime k}}),\\ T_{A^{\prime}_{1\cdots j}}(\Gamma^{\mathcal{E}^{k}_{RA^{\prime k}}})\geq 0\quad\forall j\in\{1,\ldots,k\}\end{array}\right\}$ (80)
where the map $T_{R}$ is the partial transpose map acting on system $R$.
@@
$\leq\max_{\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in\text{EXT-PPT}_{k}}\operatorname{Tr}[(\Pi_{A^{\prime}A}^{\operatorname{sym}}\otimes I_{RB})\mathcal{E}_{R\rightarrow A^{\prime}}^{k}(\psi_{RAB})]$ (83)
$=\frac{1+\widetilde{F}_{s}^{2}(\rho_{AB},k)}{2},$ (84)
where EXT-PPTk denotes the set of channels satisfying the constraints in (80). Consider the following bound for a $k$-extendible state $\omega_{AB}^{k}$ [68, Theorem II.7’] (see also just above [68, Theorem II.2] for their norm convention):
@@
We can use it and the result of [86, Lemma 7] to conclude that
@@
Then consider that, for every fixed choice of $\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in$ EXT-PPTk, there exists an entanglement-breaking channel $\mathcal{E}$ satisfying
@@
Since the inequality holds for every $\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in$ EXT-PPTk, it follows that
$\max_{\mathcal{E}_{R\rightarrow A^{\prime}}^{k}\in\text{EXT-PPT}_{k}}\operatorname{Tr}[(\Pi_{A^{\prime}A}^{\operatorname{sym}}\otimes I_{RB})\mathcal{E}_{R\rightarrow A^{\prime}}^{k}(\psi_{RAB})]\leq\frac{1+F_{s}(\rho_{AB})}{2}+\frac{2\left|R\right|\left|A^{\prime}\right|^{2}}{k}.$ (91)
@@ Appendix G Further Simulations and Details
The input states and parameterized unitaries were generated using the hardware efficient ansatz (HEA) [87] for all the simulations in our work. The HEA consists of several layers, each composed of two parameters per qubit per layer, specifying rotations about the $x$- and $y$-axes. After each layer of rotations is a series of neighboring qubit CNOT gates. When using the HEA to generate the input states, we keep the rotation angles fixed, thus leading to a fixed input state. For the parameterized unitaries, the rotation angles are parameters and are optimized over.
In Figure 10(a), we report simulation results after generating a random bipartite product state, with each partition containing two qubits. We remove all the CNOT gates from the HEA that generates the input state $\rho$ to guarantee a product state. We calculated the fidelity of separability using both the local reward function of the VQSA and the benchmark $\widetilde{F}^{1}_{s}$, the latter discussed in Appendix E.
@@
We generated all parameterized unitary circuits in the following fashion. We used the Qiskit Aer simulator and Qiskit’s Simultaneous Perturbation Stochastic Approximation (SPSA) optimizer to perform the classical optimization. All other details can be found in Table 1. The local reward function of the VQSA requires more classical processing (like picking a qubit at random to measure) and seems to require more iterations to reach the correct value. However, these downsides are outweighed by the fact that it is less susceptible to the emergence of barren plateaus. More details about the local cost function can be found in Appendix L.
@@ Appendix I Multipartite Scenarios
This appendix discusses a multipartite generalization of the separability tests for mixed states.
@@
$\rho_{A_{1}\cdots A_{M}}=\sum_{x\in\mathcal{X}}p(x)\psi^{x,1}_{A_{1}}\otimes\cdots\otimes\psi^{x,M}_{A_{M}},$ (92)
where $\{p(x)\}_{x\in\mathcal{X}}$ is a probability distribution and $\psi_{A_{i}}^{x,i}$ is a pure state for all $x\in\mathcal{X}$ and $i\in\{1,\ldots,M\}$.
@@
$\max_{\sigma_{A_{1}\cdots A_{M}}\in M\text{-}\operatorname{SEP}}F(\rho_{A_{1}\cdots A_{M}},\sigma_{A_{1}\cdots A_{M}})=\max_{\{(q(x),\varphi_{A_{1}\cdots A_{M}}^{x})\}_{x}}\sum_{x}q(x)F_{s}(\varphi^{x}_{A_{1}\cdots A_{M}}),$ (93)
@@
$\max_{\mathcal{E}\in\operatorname{EB}_{M-1}}\operatorname{Tr}[\Pi_{(A^{M-1})^{\prime}(A^{M-1})}^{\operatorname{sym}}\mathcal{E}_{R\rightarrow A^{\prime}_{1}\cdots A^{\prime}_{M-1}}(\psi_{RA^{M}})]=\frac{1}{2}\left(1+\max_{\sigma_{A_{1}\cdots A_{M}}\in M\text{-}\operatorname{SEP}}F(\rho_{A_{1}\cdots A_{M}},\sigma_{A_{1}\cdots A_{M}})\right).$ (97)
@@
where $\{\mu^{x}_{R}\}_{x}$ is a rank-one POVM and $\{\phi^{x,i}_{A^{\prime}_{i}}\}_{x,i}$ is a set of pure states. The prover sends the systems $(A^{M-1})^{\prime}=A^{\prime}_{1}\cdots A^{\prime}_{M-1}$ to the verifier. Now, the verifier performs a collective swap test on $A_{1}\cdots A_{M}$, as depicted in the final part of the circuit diagram in Figure 11. The acceptance probability of this interactive proof system is thus given by
@@
is the projector onto the symmetric subspace of $A^{\prime}$ and $A$ and $F_{(A_{1}\cdots A_{M-1})^{\prime}A_{1}\cdots A_{M-1}}$ is a tensor product of individual swaps $F_{A^{\prime}_{i}A_{i}}$, for $i\in\{1,\ldots,M-1\}$. That is,
$F_{(A_{1}\cdots A_{M-1})^{\prime}A_{1}\cdots A_{M-1}}=\bigotimes_{i=1}^{M-1}F_{A^{\prime}_{i}A_{i}}.$ (101)
@@
where $\mathcal{V}_{R^{\prime}S\rightarrow DG}$ is the unitary channel corresponding to the unitary operator $V_{R^{\prime}S\rightarrow DG}$. By reasoning similar to that in (25), (32), and (33), we find that
@@
where we have used the shorthand $\mathcal{V}\equiv\mathcal{V}_{R^{\prime}S\rightarrow DG}$. Consider that, for all $x$,
@@
$\mathcal{W}_{G\rightarrow R^{\prime}S}(\cdot)\coloneqq(V_{R^{\prime}S\rightarrow DG})^{{\dagger}}(|1\rangle\!\langle 1|_{D}\otimes(\cdot)_{G})V_{R^{\prime}S\rightarrow DG},$ (124)
and the corresponding isometry $W_{G\to R^{\prime}S}$ as $W_{G\to R^{\prime}S}\coloneqq(V_{R^{\prime}S\rightarrow DG})^{{\dagger}}|1\rangle_{D}$. Then, the acceptance probability is given by
@@
Observe that the map $\rho_{AB}\mapsto\Pi^{\operatorname{sym}}_{AA^{\prime}}(\rho_{AB}\otimes I_{A^{\prime}})\Pi^{\operatorname{sym}}_{AA^{\prime}}$ is proportional to that used in a $1\to 2$ universal cloning machine [94, Eq. (17)]. If $\rho_{AB}$ is pure, so that we write it as $\psi_{AB}$, then the following inequality holds:
@@
Quantum statistical zero-knowledge (QSZK) consists of all problems that can be solved by the interaction between a quantum verifier and a quantum prover, such that the verifier accumulates statistical evidence about the answer to a decision, but does not learn anything other than the answer by interacting with the prover [95, 56]. A complete problem for this class is quantum state distinguishability, in which the goal is to decide whether two states $\rho_{0}$ and $\rho_{1}$, generated by quantum circuits, are far or close in trace distance [95]. This is a nice problem for understanding the basics of the QSZK complexity class: the interaction begins with the verifier picking one of the states uniformly at random, recording the choice as a bit $x$, and then sending the chosen state $\rho_{x}$ to the prover over a quantum channel. The prover can then perform the optimal Helstrom measurement [96, 97] to distinguish the states, which has success probability equal to
@@
Let us define the local and global reward functions. Let $Z_{i}$ be the event of measuring zero in the $i$-th register. We then set the local reward function to be the probability of measuring zero in a register chosen uniformly at random; that is, it is given by the following:
@@
We are interested in determining inequalities related to the global and local reward functions, and the following analysis employs the same ideas used in [82, Appendix C]. Using DeMorgan’s laws, we find that
@@
$G=\Pr\!\left(\bigcap_{i}Z_{i}\right)$ $\leq 1-\frac{1}{n}\sum_{i}\left(1-\Pr\!\left(Z_{i}\right)\right)$ (174)
$=\frac{1}{n}\sum_{i}\Pr\!\left(Z_{i}\right)=L.$ (175)